OpenStudy (anonymous):

The hypotenuse of a right triangle is 20 centimeters, and the shorter leg is 12 centimeters. Find the length of the other leg.

OpenStudy (mertsj):

Try the Pythagorean Theorem

OpenStudy (anonymous):

20^2 - 12^2 = B^2 400-144= 256 sqrt of 256 = 16

OpenStudy (anonymous):

16 cm \[2\sqrt{2}cm\] \[4\sqrt{34}cm\]